A loaded sled weighing 1,250 lb is given a speed of 25.0 mi/h while 

moving a distance of 140 ft from rest on a horizontal ice surface. If the coefficient of friction 

is 0.105, what constant force, applied horizontally, would be necessary to produce this 

motion?



1
Expert's answer
2021-04-05T10:34:03-0400

The net force, acting on the sled is "F - F_f", where "F" is the applied force, and "F_f = \mu m g" is the force of friction ("\mu" is the coefficient of friction).

Hence, according to 2nd Newton's law, "m a = F - F_f = F - \mu m g", from where "F = m (a + \mu g)".

The distance for an accelerated motion can be calculated as "s = \frac{v^2-v_0^2}{2 a}"(here "v_0 = 0" is the initial speed and "v = 25 \frac{mi}{h}" is the speed after covering the distance "s = 140 ft"). From the last equation, acceleration is "a = \frac{v^2}{2 s}".

Substituting the acceleration into the expression for the force, obtain:

"F = m \left(\frac{v^2}{2 s} + \mu g\right) \approx 1413 N".